# Deep learning empowered sensor fusion to improve infant movement classification

### v1

@@ Abstract
There is a recent boom in the development of AI solutions to facilitate and enhance diagnostic procedures for established clinical tools. To assess the integrity of the developing nervous system, the Prechtl general movement assessment (GMA) is recognized for its clinical value in the diagnosis of neurological impairments in early infancy. GMA has been increasingly augmented through machine learning approaches intending to scale-up its application, circumvent costs in the training of human assessors and further standardize classification of spontaneous motor patterns. Available deep learning tools, all of which are based on single sensor modalities, are however still considerably inferior to that of well-trained human assessors. These approaches are hardly comparable as all models are designed, trained and evaluated on proprietary/ silo-data sets. We propose a sensor fusion approach for assessing fidgety movements (FMs) comparing three different sensor modalities (pressure, inertial, and visual sensors). Various combinations and two sensor fusion approaches (late and early fusion) for infant movement classification were tested to evaluate whether a multi-sensor system outperforms single modality assessments. The performance of the three-sensor fusion (classification accuracy of 94.5%) was significantly higher than that of any single modality evaluated, suggesting the sensor fusion approach is a promising avenue for automated classification of infant motor patterns. The development of a robust sensor fusion system may significantly enhance AI-based early recognition of neurofunctions, ultimately facilitating early implementation of automated detection of neurodevelopmental conditions.
@@ 1 Introduction
Different sensor modalities for AI-driven GMA have been developed, each with its own strengths and limitations [10, 1, 2, 3, 4, 5, 6, 7]. Most of these automated approaches were developed based on visual data (with RGB/RGB-D imaging sensors; [11, 12, 13, 7, 14, 15, 16, 17, 18, 19, 20, 21]), which is also the data source of ‘original GMA’. Methods using marker-based body tracking [22] or wearable sensors, e.g., inertial measurement units (IMUs) [23, 24, 25, 26, 27, 28, 29, 30], encouraged by their success in adult motion tracking, were also proposed for use in infants. Pressure sensing devices have been employed in the assessment of general movements [5, 31, 32, 10], since they are non-intrusive and easy to use, hence being particularly suitable for clinical practice. At present, the development of automated GMA approaches is still in its infancy, mostly based on small datasets, with minimal data-sharing or pooling, and targeting only a fraction or a specific aspect of the tasks involved in the standard GMA. Most AI methods focus understandably on the classification of fidgety movements (e.g., present or absent, typical or atypical; [7, 10, 18, 1, 2, 3, 4, 5, 6], a general movements pattern of high diagnostic value for the early detection of neurodevelopmental integrity at around 3 months postterm age [8, 9, 33]. The classification performances of single modality methods, irrespective of their task specifics and different participants samples, compete with each other with an accuracy level of about 90%. This is considerably inferior to that of well-trained human assessors while the methodologies cannot be compared to each other as all models are designed, trained and evaluated on proprietary data sets [1, 2, 3, 4, 5, 6, 10, 34, 18, 7, 27, 35, 36].
@@
Targeting a higher accuracy, we proposed a multi-sensory recording setup and sensor fusion approach for automated GMA, utilizing different combinations of video cameras, pressure sensing devices, and IMUs [37, 5, 26, 27, 35, 36]. The rationale behind the sensor fusion approach is that each motion tracking modality captures different types and dimensions of movement information (e.g., position and amplitude in space, body parts involved, force, velocity, frequency, direction, angular velocity and acceleration, etc.) which other sensors may miss or are unable to track directly. If different inputs are augmented, they may compensate and boost each other towards better performance. This approach needs to be empirically tested. To the best of our knowledge, there exists no public-accessible multi-sensory dataset of general movements, while no study has carried out any empirical comparisons of utilities of different sensing modalities, let alone their combinations, for analyzing infant motor functions and enhancing classification accuracy.
@@
To fill in this knowledge gap, we generated a labeled and fully synchronized multi-sensory dataset of infant movements. With this dataset, we developed a sensor fusion approach and undertook a deep learning empowered comparison of three different sensor modalities, pressure, inertial, and visual sensors and their various combinations for movement classification to evaluate if such a system outperforms single modality assessments.
@@ 2 Methods
We performed classification experiments using convolutional neural network architectures (CNN) and compared two different sensor fusion approaches, early (using one neural network) vs. late (using output of multiple neural networks) sensor fusion to address three research questions: (1) Do performances of different sensor modalities differ from each other for the same task (i.e., tracking and classifying fidgety vs. non-fidgety movements)? (2) Does sensor fusion outperform single modality assessments and lead to higher accuracy in infant movement classification? (3) Is a sensor fusion approach with non-intrusive sensors sufficient for accurate movement tracking and classification?
@@ 2.1 Method overview
Figure 1: Flow diagram of the study pipeline. N corresponds to the number of snippets (5 $s$ data units) in each step. T1-T7 correspond to seven recording sessions in biweekly intervals, starting at 4 weeks post-term age. FM- and FM+ corresponds to the presence or absence of fidgety movements. VID – video data, MAT – pressure mat data, IMU – inertial measurement unit data.
In this study, we adopted movement data captured by three different sensor modalities: 2D RGB cameras, a pressure sensing mat, and six inertial measurement units (IMUs). For movement classification based on video data, we used skeleton features as presented in prior studies [18, 34]. For movement classification based on pressure mat data, we used center of pressure (CoP) features as presented in a prior study [10]. Data from the IMUs and IMU-based classification are reported here for the first time. Based on our previous work, for the movement classification we used convolutional neural network (CNN) architectures, which is superior for the current task compared to other classification methods such as multi-layer perceptron (MLP), support vector machine (SVM) and long short-term memory (LSTM) network [10].
@@ 2.2.1 Participants
Fifty-one infants born between 2015 and 2017 from monolingual German speaking families were sampled. All parents of the participating infants provided written informed consent to study participation and publication of depersonalized data. Infant inclusion criteria were: uneventful pregnancy, uneventful delivery at term age ($>37$ weeks of gestation), singleton birth, appropriate birth weight, uneventful neonatal period, inconspicuous hearing and visual development. All parents completed high-school or higher level of education and had no record of alcohol or substance abuse.
@@ 2.2.2 Multi-modal movement recordings
According to the GMA manual [38], 5 to 8 weeks of post-term age mark a period (periods T2 and T3) of transitional movmentes which is considered a “grey”-zone between the writhing and the fidgety movement (FM) periods, which is not ideal for assessing infant general movements. FMs are most pronounced in typically developing infants from 12 weeks of post-term age on-wards (corresponding to the T5-T7 periods) [38]. Therefore, to analyze infant general movements, recordings from T1 as the “pre-fidgety period” and T5-T7 as the “fidgety period” were used. Each session, infants were dressed with specifically designed body-suits (see below) and placed in a supine position in a standard crib by the parent. We recorded infant movements using an RGB camera, a pressure sensing mat, and six inertial measurement units (IMU). All sensor recordings were synchronized [35].
For video recordings, we used a standard HD camcorder with a resolution of 1920$\times$1080 pixels, and 50 frames per second [18]. The pressure data was acquired using a Conformat pressure sensing mat (Tekscan, Inc., South Boston, Massachusetts, USA [39]). The mat was laid on the crib mattress and covered by a cotton sheet. The pressure sensing mat contains 1024 pressure sensors arranged in a 32$\times$32 grid array on an area of 471.4$\times$471.4 $mm^{2}$. The pressure mat produced pressure image frames (see Figure 2a) with a resolution of 32$\times$32 pixels, and 100 $Hz$ sampling rate [10]. To capture infant motion using wearable sensors, six wireless Xsense MTw Awinda IMUs [40] were applied. Four were each fed into a designated pocket of a customized body-suit and attached to the infant’s shoulders and the outsides of the hips to assess proximal movement features. The other two IMUs were each put into a sock and attached to the soles of the infant’s feet for distal features in the lower extremities (for sensor locations see Figure 2b). Each IMU sensor contains a 3-axis accelerometer and a 3-axis gyrometer, which measure acceleration and angular velocity in $X$, $Y$ , and $Z$ directions with a sampling rate of 60 $Hz$.
@@ 2.2.3 Movement annotations
For this study, to train and test classifiers, we used human-annotation data (Figure 1). Annotations were performed on video recordings, which were synchronized with IMU sensors and the pressure sensor such that the same labels were used for all sensor modalities. To annotate video data, videos were first cut such that infants were overall awake and active, and not fussy. We determined the length of video snippet to be 5 seconds $s$, as a minimum length of the video for human assessors feeling confident to judge whether the fidgety movement is present (FM+) or absent (FM-) [18, 10].
@@ 2.3 Data pre-processing and feature extraction
Figure 2: Flow diagrams of the feature extraction procedures for three sensor modalities: a) pressure mat data, b) IMU sensor data, and c) video data.
@@ 2.7 Evaluation procedure and quantification measures
Table 1: Data split for the hyperparameter tuning and 9-fold cross-validation. Whole dataset consists of 1683 snippets (740 FM- and 943 FM+) obtained from 45 infants. 337 snippets (obtained from 9 infants) were used for the hyperparameter tuning and 1246 snippets (obtained from 36 infants) were used for the cross-validation.
@@
In total we trained and tested 10 best models for each single sensor modality (see Supplementary Table 2) and the combination of all sensor modalities (see Supplementary Table 3). For training, we split the training set into training (1/8 of the training data) and validation (7/8 of the training data) subsets. For each fold we trained each model 20 times and then selected the model with the lowest loss score on the validation set, which was then evaluated on the test set.
For the comparison of the classification performances, we used three common classification performance measures: sensitivity (true positive rate [TPR]), specificity (true negative rate [TNR]) and balanced accuracy (BA):
@@
To compare classification accuracies of the classification models, we calculated average classification performance measures (sensitivity [$TPR$], specificity [$TNR$], and balanced accuracy [$BA$]) across nine test sets, confidence intervals of mean (CI 95%), and $p$ values for the comparison of medians using the Wilcoxon signed-rank test. Statistical significance was set at $p<0.05$.
@@ 3.1 Classification performance using single sensor modalities
Figure 4: Comparison of the classification performance of the best models of different sensor modalities on the test sets. Average balanced classification accuracy obtained from 9-fold cross-validation is shown for each case. Error bars denote confidence intervals of mean (CI $95\%$). Circles denote classification accuracies for each test set.
@@
Table 3: $p$-values for the pairwise comparisons of the classification accuracies (balanced accuracy) of different models obtained from the Wilcoxon signed-rank test.
@@ 4 Discussion
Following our experience in the use of mono-sensor approaches in the study of infant motor functions, we now propose a multimodal approach for movement recognition and classification. Using a fully synchronized multi-sensor dataset, in the present study we empirically tested and compared the utilities of three sensor modalities (pressure sensors, MAT; inertial sensors, IMU; and visual sensors, VID) and their various combinations (two- and three-sensor fusion approaches) for the same task. The task was to track and classify age-specific general movement patterns and differentiate between the absence or presence of a specific motor pattern during the first months of life, non-fidgety (FM-) vs. fidgety (FM+) movements, e.g., [47, 8, 9, 35, 34, 33].
@@
The second methodological research question was whether sensor fusion improves classification accuracy. With the sensors available at the time of study conduct and used specifically for this task, the performance of the three-sensor fusion (94.5%) was significantly more accurate than any of the single modalities. It was also significantly superior than the two-modality fusions of MAT+IMU (90.7%) or MAT+VID (91.4%). The average accuracy of the three-sensor fusion approach was again higher than the two-fold fusion of VID+IMU (93.4%), although the difference was not significant ($p=0.22$). The two-sensor fusion approaches with MAT (MAT+VID and MAT+IMU) were both significantly better than the performance of MAT alone; and the fusion with VID+IMU outperformed both MAT or IMU alone. With both two- and three-sensor fusions, the performances were superior than those of their single components, although some values were not statistically significant (Tables 2 and 3).
In our previous work, we extensively discussed possible reasons why MAT delivered moderate classification accuracy in our experiments, and the pros and cons of using divergent sensor modalities for infant motion tracking and classification [10]. With rapid technology advancement, the performance of each single modality (e.g., MAT) is likely to improve, which may further increase the accuracy of combined modalities (sensor fusion). This, however, needs to be tested in the future. For any specific task and with any single modality, the classification accuracy will eventually reach its limit (either ceiling or bottle neck). It is critical to examine, whether an approach with sensor fusion might be unnecessary (by perfect or ceiling accuracy of a single modality) or whether it is beneficial (by bottle neck of a single modality). If different dimensions of input (from diverse modalities) for the same phenomenon (e.g., motion) are augmented, the representation of the ground truth (e.g., distinguishable patterns) may be improved, hence increasing the the overall classification accuracy, as shown in our current study. If, however, different modalities only provide redundant and highly correlated information in similar dimensions, then the combined approach may not outperform its single components.

### v2

@@ Abstract
There is a recent boom in the development of AI solutions to facilitate and enhance diagnostic procedures for established clinical tools. To assess the integrity of the developing nervous system, the Prechtl general movement assessment (GMA) is recognized for its clinical value in diagnosing neurological impairments in early infancy. GMA has been increasingly augmented through machine learning approaches intending to scale-up its application, circumvent costs in the training of human assessors and further standardize classification of spontaneous motor patterns. Available deep learning tools, all of which are based on single sensor modalities, are however still considerably inferior to that of well-trained human assessors. These approaches are hardly comparable as all models are designed, trained and evaluated on proprietary/silo-data sets. With this study we propose a sensor fusion approach for assessing fidgety movements (FMs) comparing three different sensor modalities (pressure, inertial, and visual sensors). Various combinations and two sensor fusion approaches (late and early fusion) for infant movement classification were tested to evaluate whether a multi-sensor system outperforms single modality assessments. The performance of the three-sensor fusion (classification accuracy of 94.5%) was significantly higher than that of any single modality evaluated, suggesting the sensor fusion approach is a promising avenue for automated classification of infant motor patterns. The development of a robust sensor fusion system may significantly enhance AI-based early recognition of neurofunctions, ultimately facilitating automated early detection of neurodevelopmental conditions.
@@ 1 Introduction
Different sensor modalities for AI-driven GMA have been developed, each with its own strengths and limitations [10, 1, 2, 3, 4, 5, 6, 7]. Most of these automated approaches were developed based on visual data (with RGB/RGB-D imaging sensors; [11, 12, 13, 7, 14, 15, 16, 17, 18, 19, 20, 21]), which is also the data source of ‘original GMA’. Methods using marker-based body tracking [22] or wearable sensors, e.g., inertial measurement units (IMUs) [23, 24, 25, 26, 27, 28, 29, 30], encouraged by their success in adult motion tracking, were also proposed for use in infants. Pressure sensing devices have been employed in the assessment of general movements [5, 31, 32, 10], since they are non-intrusive and easy to use, hence being particularly suitable for clinical practice. At present, the development of automated GMA approaches is still in its infancy, mostly based on small datasets, with minimal data-sharing or pooling, and targeting only a fraction or a specific aspect of the tasks involved in the standard GMA. Most AI methods focus understandably on the classification of fidgety movements (e.g., present or absent, typical or atypical; [7, 10, 18, 1, 2, 3, 4, 5, 6], a general movements pattern of high diagnostic value for the early detection of neurodevelopmental integrity at around three months post-term age [8, 9, 33]. The classification performances of single modality methods, irrespective of their task specifics and different participants samples, compete with each other with an accuracy level of about 90%. This is considerably inferior to that of well-trained human assessors and the methodologies cannot be compared to each other as all models are designed, trained and evaluated on proprietary datasets [1, 2, 3, 4, 5, 6, 10, 34, 18, 7, 27, 35, 36].
@@
Targeting a higher accuracy, we proposed a multi-sensory recording setup and sensor fusion approach for automated GMA, utilizing different combinations of video cameras, pressure sensing devices, and IMUs [37, 5, 26, 27, 35, 36]. The rationale behind the sensor fusion approach is that each motion tracking modality captures different types and dimensions of movement information (e.g., position and amplitude in space, body parts involved, force, velocity, frequency, direction, angular velocity and acceleration, etc.) which other sensors may miss or are unable to track directly. If different inputs are inetgrated, they may compensate and boost each other towards better performance. This approach needs to be empirically tested. To the best of our knowledge, there exists no public-accessible multi-sensory dataset of general movements, while no study has carried out any empirical comparisons of utilities of different sensing modalities, let alone their combinations, for analyzing infant motor functions and enhancing classification accuracy.
@@
To fill in this knowledge gap, we present a labeled and fully synchronized multi-sensory dataset of infant movements. With this dataset, we develop a sensor fusion approach and undertake a deep learning empowered comparison of three different sensor modalities, pressure, inertial, and visual sensors and their various combinations for movement classification to evaluate if such a system outperforms single modality assessments.
We perform classification experiments using convolutional neural network (CNN) architectures and compare two different sensor fusion approaches, early (using one neural network) vs. late (combining outputs of multiple neural networks) sensor fusion to address three research questions: (1) Do performances of different sensor modalities differ from each other for the same task (i.e., tracking and classifying fidgety vs. non-fidgety movements)? (2) Does sensor fusion outperform single modality assessments and lead to higher accuracy in infant movement classification? (3) Is a sensor fusion approach with non-intrusive sensors sufficient for accurate movement tracking and classification?
@@ 2.1 Method overview
Figure 1: Flow diagram of the study pipeline. N corresponds to the number of snippets (5 $s$ data units) in each step. T1-T7 correspond to seven recording sessions in biweekly intervals, starting at 4 weeks post-term age. FM- and FM+ corresponds to the absence and presence of fidgety movements, respectively. VID – video data, MAT – pressure mat data, IMU – inertial measurement unit data.
In this study, we adopted movement data captured by three different sensor modalities: a 2D RGB camera, a pressure sensing mat, and six inertial measurement units (IMUs). For movement classification based on video data, we used skeleton features as presented in prior studies [18, 34]. For movement classification based on pressure mat data, we used center of pressure (CoP) features as presented in a prior study [10]. Data from the IMUs and IMU-based classification are reported here for the first time. Based on our previous work, for the movement classification we used convolutional neural network (CNN) architectures, which is superior for the current task compared to other classification methods such as multi-layer perceptron (MLP), support vector machine (SVM) and long short-term memory (LSTM) network [10].
@@ 2.2.1 Participants
51 infants born between 2015 and 2017 from monolingual German speaking families were sampled. All parents of the participating infants provided written informed consent to study participation and publication of depersonalized data. Infant inclusion criteria were: uneventful pregnancy, uneventful delivery at term age ($>37$ weeks of gestation), singleton birth, appropriate birth weight, uneventful neonatal period, inconspicuous hearing and visual development. All parents completed high-school or higher level of education and had no record of alcohol or substance abuse.
@@ 2.2.2 Multi-modal movement recordings
According to the GMA manual [38], 5 to 8 weeks of post-term age mark a period (periods T2 and T3) of transitional movements which is considered a “grey”-zone between the writhing and the fidgety movement (FM) periods, which is not ideal for assessing infant general movements. FMs are most pronounced in typically developing infants from 12 weeks of post-term age on-wards (corresponding to the T5-T7 periods) [38]. Therefore, to analyze infant general movements, recordings from T1 as the “pre-fidgety period” and T5-T7 as the “fidgety period” were used. Each session, infants were dressed with specifically designed body-suits (see Figure 2c) and placed in a supine position in a standard crib by the parent. We recorded infant movements using an RGB camera, a pressure sensing mat, and six inertial measurement units (IMU). All sensor recordings were synchronized [35].
For video recordings, we used a standard HD camcorder with a resolution of 1920$\times$1080 pixels, and a frame rate of 50 frames per second [18]. The pressure data was acquired using a Conformat pressure sensing mat (Tekscan, Inc., South Boston, Massachusetts, USA [39]). This pressure sensing mat contains 1024 pressure sensors arranged in a 32$\times$32 grid array on an area of 471.4$\times$471.4 $mm^{2}$. The mat was laid on the crib mattress and covered by a cotton sheet. The pressure mat produced pressure image frames (see Figure 2b) with a resolution of 32$\times$32 pixels, and 100 $Hz$ sampling rate [10]. To capture infant motion using wearable sensors, six wireless Xsense MTw Awinda IMUs [40] were applied. Each IMU sensor contains a 3-axis accelerometer and a 3-axis gyrometer, which measure acceleration and angular velocity in $X$, $Y$ , and $Z$ directions with a sampling rate of 60 $Hz$. Four IMU sensors were each fed into a designated pocket of a customized body-suit and attached to the infant’s shoulders and the outsides of the hips to assess proximal movement features. The other two IMUs were each put into a sock and attached to the soles of the infant’s feet for distal features in the lower extremities (for sensor locations see Figure 2c).
@@ 2.2.3 Movement annotations
For this study, to train and test classifiers, we used human-annotation data (see Figure 1). Annotations were performed on video recordings, which were synchronized with IMU sensors and the pressure sensor such that the same labels were used for all sensor modalities. To annotate video data, videos were first cut such that infants were overall awake and active, and not fussy. We determined the length of video snippet to be 5 seconds, as a minimum length of the video for human assessors feeling confident to judge whether the fidgety movement is present (FM+) or absent (FM-) [18, 10].
@@ 2.3 Data pre-processing and feature extraction
Figure 2: Flow diagrams of the feature extraction procedures for three sensor modalities: a) video data, b) pressure mat data, and c) IMU sensor data.
@@ 2.7 Evaluation procedure and quantification measures
Table 1: Data split for the hyperparameter tuning and 9-fold cross-validation. Whole dataset consists of 1683 snippets (740 FM- and 943 FM+) obtained from 45 infants. 337 snippets (obtained from 9 infants) were used for the hyperparameter tuning and 1346 snippets (obtained from 36 infants) were used for the cross-validation.
@@
In total we trained and tested 10 best models for each single sensor modality (see Supplementary Table 2) and the combination of all sensor modalities (see Supplementary Table 3). For training, we split the training set into training (7/8 of the training data) and validation (1/8 of the training data) subsets. For each fold we trained each model 20 times (with random initial conditions) and then selected the model with the lowest loss score on the validation set, which was then evaluated on the test set.
For the comparison of the classification performances, we used three common classification performance measures: sensitivity (true positive rate – TPR), specificity (true negative rate – TNR) and balanced accuracy – BA:
@@
To compare classification accuracies of the classification models, we calculated average classification performance measures (sensitivity [$TPR$], specificity [$TNR$], and balanced accuracy [$BA$]) across nine test sets, confidence intervals of mean (CI 95%), and $p$ values for the comparison of medians using the Wilcoxon two-sided signed-rank test. Statistical significance was set at $p<0.05$.
@@ 3.1 Classification performance using single sensor modalities
Figure 4: Comparison of classification performances of the best models of different sensor modalities on the test sets. Average balanced classification accuracy obtained from 9-fold cross-validation is shown for each case. Error bars denote confidence intervals of mean (CI $95\%$). Circles denote classification accuracies for each test set.
@@
Table 3: $p$-values for the pairwise comparisons of the classification accuracies (balanced accuracy) of different models obtained from the Wilcoxon two-sided signed-rank test.
@@ 4 Discussion
Following our experience in the use of mono-sensor approaches in the study of infant motor functions, we now propose a multimodal approach for movement recognition and classification. Using a fully synchronized multi-sensor dataset, in the present study we empirically tested and compared the utilities of three sensor modalities (pressure sensor, MAT; inertial sensors, IMU; and visual sensor, VID) and their various combinations (two- and three-sensor fusion approaches) for the same task. The task was to track and classify age-specific general movement patterns and differentiate between the absence or presence of a specific motor pattern during the first months of life, non-fidgety (FM-) vs. fidgety (FM+) movements, e.g., [47, 8, 9, 35, 34, 33].
@@
The second methodological research question was whether sensor fusion improves classification accuracy. With the sensors available at the time of study conduct and used specifically for this task, the performance of the three-sensor fusion (94.5%) was significantly more accurate than any of the single modalities. It was also significantly superior than the two-sensor fusions of MAT+IMU (90.7%) or MAT+VID (91.4%). The average accuracy of the three-sensor fusion approach was again higher than the two-fold fusion of VID+IMU (93.4%), although the difference was not significant ($p=0.22$). The two-sensor fusion approaches with MAT (MAT+VID and MAT+IMU) were both significantly better than the performance of MAT alone; and the fusion with VID+IMU outperformed both MAT or IMU alone. With both two- and three-sensor fusions, the performances were superior than those of their single components, although some values were not statistically significant (Tables 2 and 3).
In our previous work, we extensively discussed possible reasons why MAT delivered moderate classification accuracy in our experiments, and the pros and cons of using divergent sensor modalities for infant motion tracking and classification [10]. With rapid technology advancement, the performance of each single modality (e.g., MAT) is likely to improve, which may further increase the accuracy of combined modalities (sensor fusion). This, however, needs to be tested in the future. For any specific task and with any single modality, the classification accuracy will eventually reach its limit (either ceiling or bottle neck). It is critical to examine, whether an approach with sensor fusion might be unnecessary (by perfect or ceiling accuracy of a single modality) or whether it is beneficial (by bottle neck of a single modality). If different dimensions of input (from diverse modalities) for the same phenomenon (e.g., motion) are augmented, the representation of the ground truth (e.g., distinguishable patterns) may be improved, hence increasing the overall classification accuracy, as shown in our current study. If, however, different modalities only provide redundant and highly correlated information in similar dimensions, then the combined approach may not outperform its single components.